# User Modeling and User Profiling: A Comprehensive Survey

### 15 Feb 2024 v1

@@ Abstract
Keywords User Modeling  $\cdot$ User Profiling  $\cdot$ User Behaviors  $\cdot$ User Preferences  $\cdot$ User Interests  $\cdot$ User Representation
@@ 1 Introduction
In the pervasive era of artificial intelligence (AI) systems, the integration of such technologies into daily life is inevitable, whether embraced consciously or not. Specifically, within the realm of widely adopted tools, information retrieval (IR) and recommender systems (RSs) adeptly and proficiently supply pertinent information to users in accordance with their information requirements, personality traits, and contextual cues. Within a context where the interaction with such systems yields a voluminous amount of personal data on a daily basis, the imperative to discern individuals’ interests, characteristics, and behaviors is fulfilled by user modeling and profiling techniques [Eke et al., 2019]. These techniques primarily aim to construct a reliable user representation (i.e., a user model or user profile) commencing from generated data [Kanoje et al., 2015]. User modeling and user profiling are pivotal in understanding user behavior and providing personalized experiences. Organizations can gain valuable insights into individual preferences and interests by analyzing user data, such as browsing history, purchase patterns, and social interactions. This, in turn, enables the delivery of tailored content, products, and services, ultimately enhancing user satisfaction and engagement.
@@
1.1 Historical overview of User Modeling and User Profiling
@@
Throughout the history of scientific literature on personalization, user modeling and user profiling fields have constantly witnessed significant advancements, as described by Kobsa [2001], Cena et al. [2019], and Eke et al. [2019], among others.
@@
Traditionally, the initial steps in this research area can be dated back to the late 1970s when the notions of “user model” and “user modeling” were introduced by the contributions of Allen, Cohen, Perrault, and Rich (i.e., [Perrault et al., 1978; Cohen and Perrault, 1979; Rich, 1979]). Their pioneering work set the stage for the following decade’s research in this field. It led to the creation of numerous application systems that gathered different types of user information with varying adaptive capabilities (e.g., [Sleeman, 1985; Allgayer et al., 1989; Wahlster and Kobsa, 1989]). In these early models, there was no distinct separation between the components used for user modeling and those used for other functions. Specifically, Perrault et al. [1978] proposed a form of modeling that aimed at understanding the beliefs and goals of a user so that they could communicate. Interestingly, dialogues were exploited to understand these goals. The paper’s approach to understanding a user’s intentions using a model of the speaker’s beliefs aligns with user modeling, where a system infers a user’s needs and preferences from their interactions. This implies that user modeling systems should tailor their responses based on the user’s profile to ensure effective communication. Cohen and Perrault [1979] highlighted the need for a nuanced representation of an agent’s mental state, including beliefs, goals, and actions, to predict and influence behavior through communication. This approach was crucial for developing the idea of conversant computers capable of recognizing and generating appropriate speech acts in interaction with users. Rich [1979] introduced one of the earliest successful user modeling techniques by exploring the vision of using stereotypes in computer systems to create models of individual users. Stereotypes, in this context, are predefined sets of characteristics that are commonly associated with certain types of users. The author discussed how these stereotypes could infer user preferences and behaviors, allowing the system to make personalized recommendations. The popular system Grundy was developed as a result of this research. It acts like a librarian, suggesting novels to users based on the stereotypes it has assigned to them. Based on user interactions, the system was designed to learn and adapt its stereotypes over time to improve its recommendations. Stereotype user modeling constituted the first attempt to differentiate a user from other users [Schiaffino and Amandi, 2009] and inspired several future works (e.g., [Ardissono and Sestero, 1995; Krulwich, 1997; Zimmerman and Kurapati, 2002]).
@@
Since the late 1980s, it became apparent that the user modeling component needed to be reusable in creating user-adaptive systems. The initial move in this direction involved the creation of generic user modeling systems. Also defined as user modeling shell systems by Kobsa [1990], a generic user modeling system (GUMS) operates as an independent user modeling component within a system during runtime, requiring developers to input it with application-specific user modeling knowledge [Kobsa, 2001]. Introduced by Finin and Drager [1986], the primary GUMS was never utilized with an application system, but it established the framework for the fundamental functionality of future general (i.e., application-independent) user modeling systems, namely providing selected user modeling services that can be configured during development. A few instances of these systems are: UMT [Brajnik and Tasso, 1994], TAGUS [Paiva and Self, 1994], BGP-MS [Kobsa and Pohl, 1994], Doppelänger [Orwant, 1994], and the um toolkit [Kay, 1995]. User modeling shells were anticipated to support intricate assumptions and sophisticated reasoning about the user, particularly in the domains where these characteristics were identified. Furthermore, they were expected to be adaptable for use in a broad spectrum of other domains.
@@
A significant turning point in user modeling and profiling research was characterized by the introduction of the expert finding and expert profiling tasks within the Enterprise Track at the Text REtrieval Conference (TREC) 2005 [Craswell et al., 2005]. This was when the field started gaining considerable attention [Eke et al., 2019]. The TREC Enterprise Track’s acknowledgment of the necessity to delve into and comprehend expertise retrieval emphasized the significance of these tasks [Balog et al., 2007]. This, in turn, paved the way for a more profound investigation and comprehension of user profiling within the realm of expert finding and expert profiling [Liang and De Rijke, 2016]. Expert finding [Balog et al., 2012] is about identifying individuals who possess knowledge or expertise in a specific area within a given dataset. The aim is to find the best matches for a user’s specific request for expertise on a topic. For instance, a user might be looking for an expert who has knowledge in a particular field, and the task would involve searching through the dataset to find individuals who have demonstrated expertise in that area. The challenge lies in accurately determining a person’s expertise based on the available data and ranking these individuals based on their relevance to the given topic. Expert profiling [Balog et al., 2012], on the other hand, is about creating a comprehensive profile of an individual’s expertise. It involves gathering all relevant information about a person’s skills and knowledge from their documents. This could include their educational background, areas of specialization, projects they have worked on, their contributions to their field, and so on. The aim is to provide a detailed overview of what the person knows and what they are capable of.
@@
The 2010s marked a significant transformation in focus towards more sophisticated user profiling methods, with a growing emphasis on personalization in various digital services, particularly in recommender systems, where researchers developed advanced algorithms to analyze user behavior and preferences for improved content recommendations. [Abel et al., 2011; Lakiotaki et al., 2011; Masthoff, 2011; Konstan and Riedl, 2012]. Innovative approaches involved the use of personality-based user adaptation, where automated methods were developed for recognizing personality traits from user behaviors [Gao et al., 2013; Gou et al., 2014; Barnett et al., 2015] and conversations [Mairesse and Walker, 2010; Ivanov et al., 2011; Wei et al., 2017]. This period also saw the proliferation of mobile devices, which led to research on user modeling in mobile and location-based contexts [Ouanaim et al., 2010; Kuflik et al., 2012]. The integration of location data was explored to provide context-aware strategies and recommendations. Context-aware user modeling gained traction as researchers aimed to understand how user preferences and behaviors change in different contexts. This included factors such as location, time, and device, leading to more adaptive and responsive systems [Adomavicius and Tuzhilin, 2011; Skillen et al., 2012; Verbert et al., 2012; Codina et al., 2015]. Concurrently, the rise of social media platforms spurred an increased interest in understanding and modeling user behavior within these online social environments [Zafarani and Liu, 2013; Gou et al., 2014; Gilbert et al., 2023]. Researchers began exploring ways to incorporate social network data into user models, applying social network analysis to understand the influence of social connections on user preferences and behaviors [Vassileva, 2012; Zhong et al., 2012; Zhou et al., 2012]. This included integrating social network data and user-generated content to create more accurate and context-aware user profiles [Piao and Breslin, 2018; Kaushal et al., 2019].
@@ 1.3 Scope and novel contributions of the survey
The scientific literature has long devoted considerable attention to the research areas of user modeling and user profiling, as illustrated in the previous sections. Throughout the years, tons of papers and numerous surveys on these topics have been published. Our survey stems primarily from a number of key issues that emerged from the study of the literature: (1) the absence of an article including a comprehensive historical overview and general taxonomy; (2) the ambiguous use of basic core terminology (i.e., “user modeling”, “user model”, “user profiling”, and “user profile”); and (3) the lack of a comprehensive and up-to-date overview of this research area since 2019, to fill the gap due to the publication date of the most recent exhaustive survey (i.e., [Eke et al., 2019]).
@@
Specifically, the novel contributions presented by our survey are listed and briefly described below:
@@
  • We provide a complete overview of the large, long-standing, and ever-growing research fields of user modeling and user profiling, both from a historical and a technical viewpoint. After retracing the most significant milestones of the literature, starting from the stereotype user modeling, introduced in 1979, to the recent studies on beyond-accuracy perspectives, a formal taxonomy is proposed, taking into account all the currently active topics in the research area, including the trends that emerged in the last few years.
@@
  • We examine the definitions associated with each key term in this research domain (i.e., “user modeling”, “user model”, “user profiling”, and “user profile”), aiming to eliminate ambiguity and confusion in their usage and proposing new, encyclopedic, and easily understandable definitions.
@@
  • We present in-depth the paradigm shifts that have occurred in recent years, especially due to technological evolution, as well as the current research directions and novel trends in the field. In particular, we illustrate and discuss the advances in the following research topics:
@@
    • Implicit vs. explicit user profiling: in recent years, studies have adopted almost solely implicit (or hybrid) approaches for data collection. True explicit profiling (i.e., user data retrieved from surveys or questionnaires) is replaced by a sort of pseudo-explicit profiling, where the explicit information is taken from public data already shared by the users for different purposes (e.g., social network accounts).
@@
    • User preferences and interests: the rise and daily use of digital platforms, such as e-commerce services and recommender systems of all kinds, have inevitably led to a steady increase in information about users’ preferences and interests. This phenomenon is obviously reflected in the context of user modeling and profiling research.
@@
    • User behavior modeling: the examination of user behaviors has evolved over time, bringing forth novel concepts, such as micro and macro behavior modeling (i.e., respectively, the immediate actions that a user takes reflecting short-term preferences, and large-scale actions that reflect a user’s long-term commitment), multi-behavior modeling (i.e., the practice of integrating diverse forms of user interactions with items, as opposed to depending on a single type), sequential behavior modeling (i.e., considering the temporal sequences of behaviors influencing user interests), hierarchical user profiling (i.e., a technique used in personalized recommender systems, particularly in e-commerce, to model users’ real-time interests at different levels of granularity), and mobile user profiling (i.e., the effort of discerning users’ interests and behavioral patterns from their mobile activities).
@@
    • User representation: several works have underlined the absence of a generalized user model, suggesting that current research tends only to concentrate on particular facets of user modeling. This led to the advent and diffusion of the concepts of universal user representation and holistic user modeling.
@@
    • Evaluation: currently, there are two main lines for evaluating the performance of a standalone user profiling approach, i.e., assess the proposed model or method based on the effectiveness of a classification task at predicting a user’s personal characteristics, or generate simulated data to aid in minimizing the volume of gathered user data while maintaining profiling efficiency and safeguarding the privacy and confidentiality of users’ personal information.
@@
    • Graph data structures: as in many other domains, also in user modeling and profiling, extensive attention has been given to exploring and adopting graph structures (including knowledge graphs), particularly in the context of online social media.
@@
    • Deep learning: the advent of deep neural network-based models has ushered in a transformative era across various domains, and user modeling and profiling have not been exceptions. We discuss contributions that employed different deep learning architectures and models, i.e., differentiable user models, attention mechanism, graph neural networks, convolutional neural networks, autoencoders, recurrent neural networks, long-short term memory, and transformers.
@@
    • Beyond-accuracy perspectives: as for deep learning architectures, the adoption of “beyond-accuracy” techniques signifies a pivotal transition in every domain, including user modeling and profiling. These approaches go beyond predictive precision, prioritizing values like privacy, fairness, and explainability. In user modeling, it ensures accurate predictions while safeguarding user privacy, addressing biases, and promoting transparency.
@@
    • Application domains: user modeling and profiling are finding novel applications in critical domains, addressing contemporary challenges. In the realm of fake news detection, these techniques analyze user behavior and content interactions to identify and combat misinformation. In social networks, even if not an unexplored domain, user profiling enhances community engagement and content recommender systems by understanding individual preferences. Cybersecurity benefits from user modeling through anomaly detection, providing a proactive approach to identifying potential threats. In Massive Open Online Course (MOOC) platforms, user modeling tailors educational content, ensuring a personalized learning experience for each student.
1.4 Structure of the survey
The presented survey is structured as outlined in the following: Section 1 introduces the article by offering a detailed historical overview of the research fields (Section 1.1) and an inspection of the literature reviews published across the decades (Section 1.2); In Section 2, we analyze the core terminology. Section 3 describes in detail the paradigm shifts and the new emerging trends, while the novel comprehensive taxonomy is proposed in Section 4. Finally, in Section 5, we discuss the findings of our survey, underlining the concepts that hold over time in the field, along with the potential emerging open challenges.

### 20 Feb 2024 v2

@@ Abstract
Keywords User Modeling  $\cdot$ User Profiling  $\cdot$ User Behavior  $\cdot$ User Preferences  $\cdot$ User Interests  $\cdot$ User Representation
@@ 1 Introduction
In the pervasive era of artificial intelligence (AI) systems, the integration of such technologies into daily life is inevitable, whether embraced consciously or not. Specifically, within the realm of widely adopted tools, information retrieval (IR) and recommender systems (RSs) proficiently supply pertinent information to users in accordance with their information requirements, personality traits, and contextual cues. Within a context where the interaction with such systems yields a voluminous amount of personal data on a daily basis, the imperative to discern individuals’ interests, characteristics, and behaviors is fulfilled by user modeling and profiling techniques [Eke et al., 2019]. These techniques primarily aim to construct a reliable user representation (i.e., a user model or user profile) commencing from generated data [Kanoje et al., 2015]. User modeling and user profiling are pivotal in understanding user behavior and providing personalized experiences. Organizations can gain valuable insights into individual preferences and interests by analyzing user data, such as browsing history, purchase patterns, and social interactions. This, in turn, enables the delivery of tailored content, products, and services, enhancing user satisfaction and engagement.
@@
1.1 Historical overview
@@
Throughout the history of scientific literature on personalization, user modeling, and user profiling fields have constantly witnessed significant advancements, as described by Kobsa [2001], Cena et al. [2019], and Eke et al. [2019], among others.
@@
Traditionally, the initial steps in this research area can be dated back to the late 1970s when the notions of “user model” and “user modeling” were introduced by the contributions of Allen, Cohen, Perrault, and Rich (i.e., [Perrault et al., 1978; Cohen and Perrault, 1979; Rich, 1979]). Their pioneering work set the stage for the following decade’s research in this field. It led to the creation of numerous application systems that gathered different types of user information with varying adaptive capabilities (e.g., [Sleeman, 1985; Allgayer et al., 1989; Wahlster and Kobsa, 1989]). In these early models, there was no distinct separation between the components used for user modeling and those used for other functions. Specifically, Perrault et al. [1978] proposed a form of modeling that aimed at understanding the beliefs and goals of a user so that they could communicate. Interestingly, dialogues were exploited to understand these goals. The paper’s approach to understanding a user’s intentions using a model of the speaker’s beliefs aligns with user modeling, where a system infers a user’s needs and preferences from their interactions. This implies that user modeling systems should tailor their responses based on the user’s profile to ensure effective communication. Cohen and Perrault [1979] highlighted the need for a nuanced representation of an agent’s mental state, including beliefs, goals, and actions, to predict and influence behavior through communication. This approach was crucial for developing the idea of conversant computers capable of recognizing and generating appropriate speech acts in interaction with users. Rich [1979] introduced one of the earliest successful user modeling techniques by exploring the vision of using stereotypes in computer systems to create models of individual users. Stereotypes, in this context, are predefined sets of characteristics that are commonly associated with certain types of users. The author discussed how these stereotypes could infer user preferences and behaviors, allowing the system to make personalized recommendations. The popular system Grundy was developed as a result of this research. It acts like a librarian, suggesting novels to users based on the stereotypes assigned to them. Based on user interactions, the system was designed to learn and adapt its stereotypes over time to improve its recommendations. Stereotype user modeling constituted the first attempt to differentiate a user from other users [Schiaffino and Amandi, 2009] and inspired several future works (e.g., [Ardissono and Sestero, 1995; Krulwich, 1997; Zimmerman and Kurapati, 2002]).
@@
Since the late 1980s, it became apparent that the user modeling component needed to be reusable in creating user-adaptive systems. The initial move in this direction involved the creation of generic user modeling systems. Also defined as user modeling shell systems by Kobsa [1990], a generic user modeling system (GUMS) operates as an independent component within a system during runtime, requiring developers to input it with application-specific user modeling knowledge [Kobsa, 2001]. Introduced by Finin and Drager [1986], the primary GUMS was never utilized with an application system, but it established the framework for the fundamental functionality of future general (i.e., application-independent) user modeling systems, namely providing selected user modeling services that can be configured during development. A few instances of these systems are: UMT [Brajnik and Tasso, 1994], TAGUS [Paiva and Self, 1994], BGP-MS [Kobsa and Pohl, 1994], Doppelänger [Orwant, 1994], and the um toolkit [Kay, 1995]. User modeling shells were anticipated to support intricate assumptions and sophisticated reasoning about the user, particularly in the domains where these characteristics were identified. Furthermore, they were expected to be adaptable for use in a broad spectrum of other domains.
@@
A significant turning point in user modeling and profiling research was characterized by the introduction of the expert finding and expert profiling tasks within the Enterprise Track at the Text REtrieval Conference (TREC) 2005 [Craswell et al., 2005]. This was when the field started gaining considerable attention [Eke et al., 2019]. The TREC Enterprise Track’s acknowledgment of the necessity to delve into and comprehend expertise retrieval emphasized the significance of these tasks [Balog et al., 2007]. This, in turn, paved the way for a more profound investigation and comprehension of user profiling within the realm of expert finding and expert profiling [Pavan et al., 2015; Liang and De Rijke, 2016]. Expert finding [Balog et al., 2012] is about identifying individuals who possess knowledge or expertise in a specific area within a given dataset. The aim is to find the best matches for a user’s specific request for expertise on a topic. For instance, a user might be looking for an expert who has knowledge in a particular field, and the task would involve searching through the dataset to find individuals who have demonstrated expertise in that area. The challenge lies in accurately determining a person’s expertise based on the available data and ranking these individuals based on their relevance to the given topic. Expert profiling [Balog et al., 2012], on the other hand, is about creating a comprehensive profile of an individual’s expertise. It involves gathering all relevant information about a person’s skills and knowledge from their documents. This could include their educational background, areas of specialization, projects they have worked on, their contributions to their field, and so on. The aim is to provide a detailed overview of what the person knows and what they are capable of.
@@
The 2010s marked a significant transformation in focus toward more sophisticated user profiling methods, with a growing emphasis on personalization in various digital services, particularly in RSs, where researchers developed advanced algorithms to analyze user behavior and preferences for improved content recommendations [Abel et al., 2011; Lakiotaki et al., 2011; Masthoff, 2011; Konstan and Riedl, 2012]. Innovative approaches involved the use of personality-based user adaptation, where automated methods were developed for recognizing personality traits from user behaviors [Gao et al., 2013; Gou et al., 2014; Barnett et al., 2015] and conversations [Mairesse and Walker, 2010; Ivanov et al., 2011; Wei et al., 2017]. A notable surge in academic literature was due to semantic user modeling techniques, which concern the creation of computational models to understand and predict user preferences, behaviors, and intentions based on semantic information derived from various data sources [Aroyo and Houben, 2010; Bakalov et al., 2010; Plumbaum et al., 2011; Cena et al., 2012; Plumbaum et al., 2012; Piao and Breslin, 2016]. This period also saw the proliferation of mobile devices, which led to research on user modeling in mobile and location-based contexts [Ouanaim et al., 2010; Kuflik et al., 2012]. The integration of location data was explored to provide context-aware strategies and recommendations. Context-aware user modeling gained traction as researchers aimed to understand how user preferences and behaviors change in different contexts. This included factors such as location, time, and device, leading to more adaptive and responsive systems [Adomavicius and Tuzhilin, 2011; Skillen et al., 2012; Verbert et al., 2012; Said et al., 2013; Codina et al., 2015]. Concurrently, the rise of social media platforms spurred an increased interest in understanding and modeling user behavior within these online social environments [Zafarani and Liu, 2013; Gou et al., 2014; Gilbert et al., 2023]. Researchers began exploring ways to incorporate social network data into user models, applying social network analysis to understand the influence of social connections on user preferences and behaviors [Vassileva, 2012; Zhong et al., 2012; Zhou et al., 2012]. This included integrating social network data and user-generated content to create more accurate and context-aware user profiles [Piao and Breslin, 2018; Kaushal et al., 2019].
@@ 1.3 Scope and novel contributions of the survey
The scientific literature has long devoted considerable attention to the research areas of user modeling and user profiling, as illustrated in the previous sections. Across the years, tons of papers and numerous surveys on these topics have been published. Our survey stems primarily from a number of key issues that emerged from the study of the literature: (1) the absence of an article including a wide historical overview and general taxonomy; (2) the ambiguous use of basic core terminology (i.e., “user modeling”, “user model”, “user profiling”, and “user profile”); and (3) the lack of a comprehensive and up-to-date overview of this research area since 2019, to fill the gap due to the publication date of the most recent exhaustive survey (i.e., [Eke et al., 2019]).
@@
Specifically, the novel contributions presented by our survey are listed and briefly described below, with the indication of the section including the related content:
@@
  • We provide a complete outline of the large, long-standing, and ever-growing research fields of user modeling and user profiling (Section 1), by offering a detailed historical overview (Section 1.1) and an inspection of the literature reviews published across the decades (Section 1.2). After retracing the most significant milestones of the literature, starting from the stereotype user modeling, introduced in 1979, to the recent studies on beyond-accuracy perspectives, a formal taxonomy is proposed (Section 4), taking into account all the currently active topics in the research area, including the trends that emerged in the last few years.
@@
  • We examine the definitions associated with each key term in this research domain (i.e., “user modeling”, “user model”, “user profiling”, and “user profile”), aiming to eliminate ambiguity and confusion in their usage and proposing new, encyclopedic, and easily understandable definitions (Section 2).
@@
  • We present in-depth the paradigm shifts that have occurred in recent years, especially due to technological evolution, as well as the current research directions and novel trends in the field (Section 3). In particular, we illustrate and discuss the advances in the following research topics:
@@
    • Implicit vs. explicit user profiling (Section 3.1): in recent years, studies have adopted almost solely implicit (or hybrid) approaches for data collection. True explicit profiling (i.e., user data retrieved from surveys or questionnaires) is replaced by a sort of pseudo-explicit profiling, where the explicit information is taken from public data already shared by the users for different purposes (e.g., social network accounts).
@@
    • User preferences and interests (Section 3.2): the rise and daily use of digital platforms, such as e-commerce services and RSs of all kinds, have inevitably led to a steady increase in information about users’ preferences and interests. This phenomenon is obviously reflected in the context of user modeling and profiling research.
@@
    • User behavior modeling (Section 3.3): the examination of user behaviors has evolved over time, bringing forth novel concepts, such as micro and macro behavior modeling (i.e., respectively, the immediate actions that a user takes reflecting short-term preferences, and large-scale actions that reflect a user’s long-term commitment), multi-behavior modeling (i.e., the practice of integrating diverse forms of user interactions with items, as opposed to depending on a single type), sequential behavior modeling (i.e., considering the temporal sequences of behaviors influencing user interests), hierarchical user profiling (i.e., a technique used in personalized recommender systems, particularly in e-commerce, to model users’ real-time interests at different levels of granularity), and mobile user profiling (i.e., the effort of discerning users’ interests and behavioral patterns from their mobile activities).
@@
    • User representation (Section 3.4): several works have underlined the absence of a generalized user model, suggesting that current research tends only to concentrate on particular facets of user modeling. This led to the advent and diffusion of the concepts of universal user representation and holistic user modeling.
@@
    • Evaluation (Section 3.5): currently, there are two main lines for evaluating the performance of a standalone user profiling approach, i.e., assess the proposed model or method based on the effectiveness of a classification task at predicting a user’s personal characteristics, or generate simulated data to aid in minimizing the volume of gathered user data while maintaining profiling efficiency and safeguarding the privacy and confidentiality of users’ personal information.
@@
    • Graph data structures (Section 3.6): as in many other domains, also in user modeling and profiling, extensive attention has been given to exploring and adopting graph structures (including knowledge graphs), particularly in the context of online social media.
@@
    • Deep learning (Section 3.7): the advent of deep neural network-based models has ushered in a transformative era across various domains, and user modeling and profiling have not been exceptions. We discuss contributions that employed different deep learning architectures and models, i.e., differentiable user models, attention mechanism, graph neural networks, convolutional neural networks, autoencoders, recurrent neural networks, long-short term memory, and transformers.
@@
    • Beyond-accuracy perspectives (Section 3.8): as for deep learning architectures, the adoption of “beyond-accuracy” techniques signifies a pivotal transition in every domain, including user modeling and profiling. These approaches go beyond predictive precision, prioritizing values like privacy, fairness, and explainability. In user modeling, it ensures accurate predictions while safeguarding user privacy, addressing biases, and promoting transparency.
@@
    • Application domains (Section 3.9): user modeling and profiling are finding novel applications in critical domains, addressing contemporary challenges. In the realm of fake news detection, these techniques analyze user behavior and content interactions to identify and combat misinformation. In social networks, even if not an unexplored domain, user profiling enhances community engagement and content recommender systems by understanding individual preferences. Cybersecurity benefits from user modeling through anomaly detection, providing a proactive approach to identifying potential threats. In Massive Open Online Course (MOOC) platforms, user modeling tailors educational content, ensuring a personalized learning experience for each student.
  • We discuss the findings of our survey, underlining the concepts that hold over time in the field, along with the emerging future research directions (Section 5).